# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 3.4.1 Segmentation
Segmentation has grown dramatically in the past few years, given its central role in visual perception. Many segmentation methods now integrate an input prompt for users to define their preferred output appearances, such as pixel-wise segmentation, bounding boxes around objects, or segmented areas of interest. Most of these methods utilize transformer architectures [149]. Among them, Segment Anything (SAM) by Meta AI [150] stands out as a pioneer in promptable segmentation approaches. This method computes masks in real-time and has been trained with over 1 billion masks across 11 million images, facilitating transferability from zero-shot to new image distributions and tasks. HQ-SAM [151] enhances SAM by incorporating global-local feature fusion, leading to high-quality mask predictions. SegGPT [152] proposed context ensemble strategies and allows users to tune a prompt for a specific dataset, scene, or even a person, while SEEM [153] provides a completely promptable and interactive segmentation interface. More recently, SAM2 [155] introduced support for real-time video segmentation. It is a unified model trained on a larger dataset than SAM. Interactive tools enable users to mark areas of interest and specify regions to exclude from the segmentation map.
@@
Applying INRs to segmentation is more popular in the medical domain, as the specific signals used, such as computed tomography (CT) and magnetic resonance imaging (MRI), can be formulated as continuous functions. In creative technologies, unsupervised domain adaptation (UDA) and INR are used for continuous rectification function modeling in [159], achieving superior segmentation results in night vision. Recently, the work integrated with a non-local means block in [270] has shown a significant improvement for instant segmentation in low-light scenes.
3D segmentation is also crucial for scene manipulation. In radiance fields, earlier segmentation methods required additional modules to separate objects from the background, such as using k-means clustering as demonstrated in [271]. However, the recent SA3D approach [160] segments 3D objects using NeRFs as the structural prior. SA3D operates by taking a trained NeRF and a set of prompts from a single view, then performing an iterative procedure. This involves rendering novel 2D views, self-prompting SAM for 2D segmentation, and projecting the segmentation back onto 3D mask grids.
@@ 3.4.2 Detection and recognition
Back in 2020, DETR by Facebook AI [161] is one of the first to adopt the transformer architecture for object detection. The approach achieves comparable results to an optimized Faster R-CNN [272], introduced in 2015. Deformable convolution has been used, Deformable DETR [162], resulting in training faster and approximately 5% accuracy improvement. A survey until 2022 reported in [273] show that Deformable DETR and Swin Transformer [16] outperform pure CNN-based YOLOv4 [274]. SwinV2 improves the first version by replacing original dot product attention with scaled cosine attention, and the accuracy is improved approximately 5%. Later, RT-DETR [275] improves inference speed by decoupling the intra-scale interaction and cross-scale fusion of features with different scales. RT-DETR is 25% faster than YOLOv8414141https://github.com/ultralytics/ultralytics with 6% improvement on MS COCO Object Detection dataset. Recently, YOLOv10 [276] has been released. YOLOv10 further improve the speed of detection approximately 30% over RT-DETR with the same accuracy. A review of transformer-based methods for object detection can be found in [277]. As estimating depth from a 2D image has been well established [183], MonoDTR [163] integrates such module to estimate 3D box of the detected object from one image. 3D object detection is also employed transformers.
While DMs are primarily used to generate synthetic datasets [278, 279], they have also been demonstrated to function as zero-shot classifiers by Li et al. [165]. DMs are of interest for detection tasks as well, though feature extractors are still predominantly based on CNNs, such as ResNet, or Transformers, like Swin. DiffusionDet [166] formulates object detection as a denoising diffusion process from noisy boxes to object boxes, reporting performance that surpasses DETR.
@@ 3.4.3 Tracking
Object tracking stands out as one of the tasks that greatly benefits from transformers as attention is needed in both space and time. An experimental survey cited in [280] reveals that transformer-based methods consistently rank at the top of the leaderboard across various datasets. In the Visual Object Tracking (VOT) challenges of 2023424242https://eu.aihub.ml/competitions/201#results, all top-10 leaders employed transformer-based methodologies. The highest-performing approach achieved a 10% improvement in tracking quality compared to the winner in 2020. The current state-of-the-art for single-object tracking [^43] https://paperswithcode.com/sota/visual-object-tracking-on-lasot, however, is based on cross-attention and Mamba [175].
The first three tracking-by-attention approaches are TrackFormer [167], MixFormer [169], and ToMP [170]. TrackFormer extracts visual features using a CNN-based encoder, which are then tracked using a vanilla transformer [10] in a frame sequence, while MixFormer introduces cross-attention between the target and search regions. ToMP tracks the objects using prediction aspects. Many more methods have been proposed, including popular ones like SeqTrack [172] and Track Anything Model (TAM) [171]. SeqTrack extracts visual features with a bidirectional transformer, while the decoder generates a sequence of bounding box values autoregressively with a causal transformer. TAM combines SAM [150] and XMem [281], offering tracking and segmentation performance on the human-selected target. However, the masked area is still not very sharp, and there is a subtle degree of temporal inconsistency. MOTRv2 [173] combines YOLOX [282] for object recognition and MOTR [168] for tracking, outperforming TrackFormer by 20%. Additionally, some methods have been specifically proposed for challenging environments, such as low light [174] and small objects, as seen in AnyFlow [179]. The later case exploits INR to upsample a continuous coordinate-based flow map, similar to SISR technique proposed in [121].
Similar to detection tasks, DMs for tracking tasks are used as downstream tasks by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal features of the videos [176]. DiffusionTrack [177] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [178] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports slightly underperforming MixFormer.
@@
Bridging the gap between digital and physical realms, 3D reconstruction and rendering are integral to various creative technologies. In film and animation, they enable the creation of detailed digital models that blend seamlessly with live-action footage. Video games leverage these technologies for dynamic environmental rendering. VR and AR use 3D reconstruction to create immersive and interactive experiences, with AR integrating digital content into real-world contexts. With recent AI technologies, 3D reconstruction and rendering have become faster and closer to reality. In particular, neural radiance fields and Gaussian Splatting enable artists and film producers to create shots that cannot be one in the real shooting environments.
@@ 3.5.1 Depth Estimation
Accurate depth information is typically required to construct 3D models. Depth sensors, such as lidar (Light Detection and Ranging) and structured-light 3D scanners, can be used for this purpose, but their applications are often limited by distance and cost. Consequently, vision-based sensors have become widely used. These sensors utilize two or more cameras to simulate human binocular vision or employ a single camera to capture images from different locations.
@@ 3.5.2 Neural Radiance Fields
Neural Radiance Fields (NeRFs), introduced in [45], have demonstrated the ability to learn a 3D scene from a smaller number of images captured from various viewpoints, as opposed to photogrammetry. They excel in neural rendering, particularly in view-dependent novel view synthesis, and have effectively tackled several challenges associated with automated 3D capture [41], such as accurately representing the reflectance properties of the scene. NeRFs offer high-resolution photo-realistic novel views and flexibility on postprocessing. They have hence gained lots of attention in cinematography [283], as it potentially safe time and cost particularly for outdoor shooting.
@@
In the NeRF process (see Fig. 7 (a)), the camera positions and orientations are typically estimated from a series of 2D images using techniques like feature-mapping and Structure-from-Motion (SfM), as demonstrated in [284]. Leveraging INR, each image (or camera pose) is mapped into camera rays that traverse the scene, generating 3D points with directional radiance (towards the camera). These points are then processed by an MLP to predict volume density and emitted radiance. Subsequently, the volume rendering techniques are employed to generate an image, which is compared with the original for loss calculation. The MLP iteratively refines the model by minimizing this loss.
@@
Since then, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [185] proposed unbounded anti-aliased technique achieving full 360 degree content. Google Research [191] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [193].
@@
While the rendering quality of NeRF is outstanding, training and rendering times remain extremely high. The Instant-NGP tool developed by Nvidia [190] enables real-time training of NeRFs by bypassing sampling in empty spaces and dense areas, and by incorporating multi-resolution hash encoding techniques. These advancements substantially reduce the computational burden associated with representing high-resolution image features – training times have been reduced from hours to just a few seconds. Moreover, it offers VR controls for immersive 3D rendering experiences using OpenXR [^44] An open-source, royalty-free standard for access to virtual reality and augmented reality platforms and devices. https://www.khronos.org/openxr/. This allows users to navigate scenes, manipulate objects, and interact with the environment directly through VR headsets. Diffusion models are integrated to regularize NeRF reconstructions [187], resulting in smoother depth continuity and clearer edges where depth discontinuities occur.
@@
The initial attempt to apply NeRF to dynamic scenes was undertaken by Pumarola et al. [189], known as D-NeRF. However, the current leading method for generating high-quality novel views of real dynamic scenes is TiNeuVox [192]. It enhances temporal information by interpolating voxel features before feeding them into the radiance network to estimate density and color, similar to ordinary NeRF. The main drawback of these methods is the large model size and/or long training time. Therefore, $K$-planes [195] propose a simple planar factorization for volumetric rendering, achieving low memory usage (1000$\times$ compression over a full 4D grid). Wavelet transform is employed in [194] to further reduce model size.
@@
3.5.3 Gaussian Splatting
The main issue with modeling NeRFs to generate high-quality novel views is the training time, which can exceed a day. 3D Gaussian Splatting (3D-GS) [196] has been introduced using anisotropic 3D Gaussians for a high-quality, unstructured representation of radiance fields. The process estimates a sparse point cloud through SfM. Each point possesses 3D Gaussian properties, such as position, covariance matrix, opacity, and spherical harmonics coefficients representing colors. The optimization of these parameters is interleaved with steps that control the density of the Gaussians to better represent the scene, as shown in Fig. 7 (b). A survey of 3D-GS can be found in [286].
@@ 3.5.3 Gaussian Splatting
For dynamic scenes, 4D Gaussian Splatting (4D-GS) [197] proposes a Gaussian deformation field for motion and shape. It exploits a multi-resolution encoding method, achieving real-time rendering of up to 82 fps at a resolution of 800$\times$800 pixels. Instead of developing in 4D, CoGS [198] exploits 3D-GS by integrating control mechanisms in separate regions to learn individual temporal dimensions. SC-GS [199] extracts sparse control points and uses an MLP to predict time-varying 6 DoF transformations. While the results show better visual quality than 4D-GS and CoGS, the performance heavily relies on camera pose estimation.
LUMA AI [^45] https://lumalabs.ai/interactive-scenes and Polycam [^46] https://poly.cam/captures offer free tools for Gaussian splat and photogrammetry creators. The 3D objects created can be experienced with VR headsets for more immersive 3D and further used or developed in other applications. However, these tools have limitations in handling dynamic scenes, which are challenging for both NeRF and GS to achieve good results.

### v3

@@ 3.4.1 Segmentation
The need for segmentation has grown dramatically in the past few years, given its central role in visual perception. Many segmentation methods now integrate an input prompt for users to define their preferred output appearances, such as pixel-wise segmentation, bounding boxes around objects, or segmented areas of interest. Most of these methods utilize transformer architectures [161]. Among them, Segment Anything (SAM) by Meta AI [162] stands out as a pioneer in promptable segmentation approaches. This method computes masks in real-time and has been trained with over 1 billion masks across 11 million images, facilitating transferability from zero-shot to new image distributions and tasks. HQ-SAM [163] enhances SAM by incorporating global-local feature fusion, leading to high-quality mask predictions. SegGPT [164] proposed context ensemble strategies and allows users to tune a prompt for a specific dataset, scene, or even a person, while SEEM [165] provides a completely promptable and interactive segmentation interface. More recently, SAM2 [167] introduced support for real-time video segmentation. It is a unified model trained on a larger dataset than SAM. Interactive tools enable users to mark areas of interest and specify regions to exclude from the segmentation map. Zhou et al. propose an audio-visual segmentation (AVS) to generate pixel-level segmentation masks for sounding objects in audible videos. DVIS++ by [168] introduces a universal video segmentation framework capable of producing instance, semantic, and panoptic segmentation outputs. This transformer-based architecture comprises a segmentor, tracker, and refinement module, achieving state-of-the-art performance across several video segmentation benchmarks.
@@
Applying INRs to segmentation is more popular in the medical domain, as the specific signals used, such as computed tomography (CT) and magnetic resonance imaging (MRI), can be formulated as continuous functions. In creative technologies, unsupervised domain adaptation (UDA) and INRs are used for continuous rectification function modeling in [172], achieving superior segmentation results in night vision. Recently, this work has been integrated with a non-local means block in [308] showning a significant improvement for instant segmentation in low-light scenes.
3D segmentation is also crucial for scene manipulation. In radiance fields, earlier segmentation methods required additional modules such as using k-means clustering to separate objects from the background [309]. However, the recent SA3D approach [173] segments 3D objects using NeRFs as the structural prior. SA3D operates by taking a trained NeRF and a set of prompts from a single view, then performing an iterative procedure. This involves rendering novel 2D views, self-prompting SAM for 2D segmentation, and projecting the segmentation back onto 3D mask grids. A comprehensive survey of 3D segmentation in computer vision can be found in [310].
@@ 3.4.2 Detection and recognition
Introduced in 2020, DETR by Facebook AI [174] was one of the first to adopt a transformer architecture for object detection. The approach achieves comparable results to an optimized Faster R-CNN [311], introduced in 2015. Deformable convolution has alson been used, (Deformable DETR [175]), resulting in training faster with approximately 5% accuracy improvement. A survey until 2022 [312] reported that Deformable DETR and Swin Transformers [16] outperform pure CNN-based YOLOv4 [313]. SwinV2 improves the first version by replacing original dot product attention with scaled cosine attention, improving accuracy by approximately 5%. Later, RT-DETR [314] improved inference speed by decoupling the intra-scale interaction and cross-scale fusion of features with different scales. RT-DETR is 25% faster than YOLOv8464646https://github.com/ultralytics/ultralytics with 6% improvement on MS COCO Object Detection dataset. Recently, YOLOv10 [315] has been released. YOLOv10 further improves the speed of detection approximately by 30% over RT-DETR with the same accuracy. A review of transformer-based methods for object detection can be found in [316, 317]. Recently, YOLO12 [180] introduced an attention-centric architecture, achieving a 2.1% and 1.2% mAP improvement over YOLOv10-N and YOLOv11-N respectively, with only a slight decrease in speed.
To detect 3D objects, the transformer-based method MonoDTR [177] incorporates depth estimation from a single 2D image [201] to predict 3D bounding boxes. More 3D object detection methods have been developed for autonomous driving [318]; however, these approaches can also be adapted for AR and VR applications [179].
While DMs are primarily used to generate synthetic datasets [319, 320], they have also been demonstrated to function as zero-shot classifiers by Li et al. [181]. DMs are also of interest for detection tasks, Although feature extractors are still predominantly based on CNNs, such as ResNet, or Transformers (like Swin). DiffusionDet [182] formulates object detection as a denoising diffusion process from noisy boxes to object boxes, reporting performance that surpasses DETR. DMs have also been employed for anomaly detection [183, 184], functioning similarly to zero-shot classifiers.
@@ 3.4.3 Tracking
Object tracking stands out as one of the tasks that greatly benefits from transformers since attention is needed in both space and time. An experimental survey cited in [321] reveals that transformer-based methods consistently rank at the top of the leaderboard across various datasets. In the Visual Object Tracking (VOT) challenges of 2023474747https://eu.aihub.ml/competitions/201#results, all of the top-10 employed transformer-based methodologies. The highest-performing approach achieved a 10% improvement in tracking quality compared to the winner in 2020. The current state-of-the-art for single-object tracking [^48] https://paperswithcode.com/sota/visual-object-tracking-on-lasot, however, is based on cross-attention and Mamba [193].
The first three tracking-by-attention approaches are TrackFormer [185], MixFormer [187], and ToMP [188]. TrackFormer extracts visual features using a CNN-based encoder, which are then tracked using a vanilla transformer [10] in a frame sequence, while MixFormer introduces cross-attention between the target and search regions. ToMP tracks the objects using prediction aspects. Many more methods have been proposed, including SeqTrack [190] and Track Anything Model (TAM) [189]. SeqTrack extracts visual features with a bidirectional transformer, while the decoder generates a sequence of bounding box values autoregressively with a causal transformer. TAM combines SAM [162] and XMem [322], offering tracking and segmentation performance on the human-selected target. However, the masked area is still not very sharp, and there is a subtle degree of temporal inconsistency. MOTRv2 [191] combines YOLOX [323] for object recognition and MOTR [186] for tracking, outperforming TrackFormer by 20%. Additionally, some methods have been specifically proposed for challenging environments, such as low light [192] and small objects, as seen in AnyFlow [197]. The latter exploits INR to upsample a continuous coordinate-based flow map, similar to SISR technique proposed in [127].
Similarly to detection tasks, DMs for tracking tasks are used as downstream processes by concatenating the diffusion head to the feature extraction backbone. However, a spatial-temporal fusion module has been added to the diffusion head to exploit temporal video features[194]. DiffusionTrack [195] localizes the target in a progressive diffusion manner, which is claimed to better handle challenging scenarios. The method in [196] exploits spatial-temporal weighting to suppress the probability of the tracker changing the target to the distractors. It, however, reports under-performance compared to MixFormer.
@@
Bridging the gap between digital and physical realms, 3D reconstruction and rendering are integral to various creative technologies. In film and animation, they enable the creation of detailed digital models that blend seamlessly with live-action footage. Video games and digital twins leverage these technologies for dynamic environmental rendering. VR and AR use 3D reconstruction to create immersive and interactive experiences, with AR integrating digital content into real-world contexts. With recent AI technologies, 3D reconstruction and rendering have become faster and closer to reality. In particular, neural radiance fields and Gaussian Splatting enable artists and film producers to create shots that cannot be one in the real shooting environments.
@@ 3.5.1 Depth Estimation
Accurate depth information (alongside texture data) is typically required to construct 3D models. Depth sensors, such as lidar (Light Detection and Ranging) and structured-light 3D scanners, can be used for this purpose, but their applications are often limited by distance and cost. Consequently, vision-based sensors have become widely used. These sensors utilize two or more cameras to simulate human binocular vision or employ a single camera to capture images from different locations.
@@ 3.5.2 Neural Radiance Fields
Neural Radiance Fields (NeRFs), introduced in [43], have demonstrated the ability to learn a 3D scene from a smaller number of images captured from various viewpoints, as opposed to photogrammetry. They excel in neural rendering, particularly in view-dependent novel view synthesis, and have effectively tackled several challenges associated with automated 3D capture [39], such as accurately representing the reflectance properties of the scene. NeRFs offer high-resolution photo-realistic novel views and flexibility in postprocessing. They have hence gained significant attention in cinematography [324], as they offer reduced time and cost, particularly for outdoor shooting.
@@
In the NeRF process (see Fig. 7 (a)), the camera positions and orientations are typically estimated from a series of 2D images using techniques like feature-mapping and Structure-from-Motion (SfM), as demonstrated in [325]. Leveraging INR, each image (or camera pose) is mapped into camera rays that traverse the scene, generating 3D points with directional radiance (towards the camera). These points are then processed by an MLP to predict volume density and emitted radiance. Subsequently, volume rendering techniques are employed to generate an image, which is compared with the original via loss calculation. The MLP iteratively refines the model by minimizing this loss.
@@
Since their introduction, there have been many variants of NeRFs aimed at improving their performance. Mip-NeRF360 [204] proposed unbounded anti-aliased technique achieving full 360 degree content. Google Research [210] trains NeRF from noisy RAW images captured in the dark scene, allowing changing viewpoint, focus, exposure, and tone mapping simultaneously. With segmentation techniques significantly advanced (see Section 3.4.1), there have been integrations utilizing semantic segmentation to enhance 3D representation [212]. DSEM-NeR [203] integrates the pretrained CLIP model to extract multimodal features—including color, depth, and semantics—from multi-view 2D images, thereby enhancing the reconstruction quality of complex scenes.
@@
While the rendering quality of NeRF is very good, training and rendering times remain extremely high. The Instant-NGP tool developed by Nvidia [209] enables real-time training of NeRFs by bypassing sampling in empty spaces and dense areas, and by incorporating multi-resolution hash encoding techniques. These advancements substantially reduce the computational burden associated with representing high-resolution image features – training times have been reduced from hours to just a few seconds. Moreover, it offers VR controls for immersive 3D rendering experiences using OpenXR [^49] An open-source, royalty-free standard for access to virtual reality and augmented reality platforms and devices. https://www.khronos.org/openxr/. This allows users to navigate scenes, manipulate objects, and interact with the environment directly through VR headsets. Diffusion models are integrated to regularize NeRF reconstructions [206], resulting in smoother depth continuity and clearer edges where depth discontinuities occur.
@@
The initial application of NeRFs to dynamic scenes was undertaken by Pumarola et al. [208], known as D-NeRF. However, the current leading method for generating high-quality novel views of real dynamic scenes is TiNeuVox [211]. It enhances temporal information by interpolating voxel features before feeding them into the radiance network to estimate density and color, similar to ordinary NeRF. DynVideo-E [214] adds an MLP to predict motion fields but focuses on human-centric content. PaReNeRF [217] addresses large-scale dynamic scenes using patch-based sampling. The main drawback of these methods is the large model size and/or long training time. Therefore, $K$-planes [218] propose a simple planar factorization for volumetric rendering, achieving low memory usage (1000$\times$ compression over a full 4D grid). Wavelet transform are employed in [215] to further reduce model size. KFD-NeRF [216] incorporates a Kalman filter-guided deformation field for more accurate motion estimation.
@@
3.5.3 3D Gaussian Splatting
The main issue with NeRFs as a method to generate high-quality novel views is training time, which can exceed a day for high-resolution content on a single RTX 3090 GPU [213]. 3D Gaussian Splatting (3D-GS) [219] has been introduced to address this, using anisotropic 3D Gaussians to form a high-quality, unstructured representation of radiance fields. The process estimates a sparse point cloud through SfM. Each point possesses 3D Gaussian properties, such as position, covariance matrix, opacity, and spherical harmonics coefficients representing colors. The optimization of these parameters is interleaved with steps that control the density of the Gaussians to better represent the scene, as shown in Fig. 7 (b). A survey of 3D-GS can be found in [327].
@@
For dynamic scenes, 4D Gaussian Splatting (4D-GS) [220] introduces a Gaussian deformation field for motion and shape. It exploits a multi-resolution encoding method, achieving real-time rendering of up to 82 fps at a resolution of 800$\times$800 pixels on an RTX 3090 GPU. Instead of developing in 4D, CoGS [221] exploits 3D-GS by integrating control mechanisms in separate regions to learn individual temporal dimensions. SC-GS [222] extracts sparse control points and uses an MLP to predict time-varying 6 DoF transformations. While the results show better visual quality than 4D-GS and CoGS, the performance heavily relies on camera pose estimation. Kong et al. [225] represent dynamic scenes using sparse, time-variant attribute modeling with a deformable MLP, while efficiently filtering out anchors corresponding to static regions. Their model achieves fast rendering speeds of over 110 FPS at a resolution of 960$\times$540—nearly 10 times faster than SC-GS—and delivers a 1 dB improvement in PSNR.
LUMA AI [^50] https://lumalabs.ai/interactive-scenes and Polycam [^51] https://poly.cam/captures offer free tools for Gaussian splatting and photogrammetry creation for non-commercial use. The 3D objects created can be experienced with VR headsets for more immersive 3D and further used or developed in other applications. However, these tools have limitations in handling dynamic scenes due to occlusions, sparse observations per timestamp, and object reappearances over time. Rendering dynamic avatars can produce higher quality results by incorporating additional information. For example, EVA [224] disentangles the 3D Gaussian appearance into skeletal motion, facial expressions, body movements, and skin. These components are then splatted to render the final photorealistic image.
3.5.4 Digital Twins
A digital twin is a virtual replica of a physical object, system, or process, continuously updated with real-time data for purposes such as simulation, testing, monitoring, and maintenance. This technology is increasingly adopted across various applications within the creative industries. For example, in product design and branding, it enables immediate observation of how a design performs in various contexts, facilitating the development of user-friendly products. Unilever reported that integrating digital product twins with 3D technologies, such as NVIDIA Omniverse [^52] https://www.nvidia.com/en-gb/omniverse/, enabled the creation of product imagery twice as fast and 50% more cost-effective [^53] https://www.unilever.com/news/press-and-media/press-releases/2025/unilever-reinvents-product-shoots-with-digital-twins-and-ai/. Digital twins also allow consumers to explore products or spaces virtually, simulating real-world interactions.
Accenture plc, a global professional services company, collaborated with Walt Disney Studios to develop digital twin technologies aimed at transforming the filmmaking process [^54] https://www.accenture.com/mx-es/case-studies/communications-media/empowering-film-creatives-digital-twins. Their goal is to generate remotely accessible 3D models, enabling virtual exploration of potential shooting locations without requiring physical visits. The Virtual StudioLAB provides a digital replica created using 360-degree imagery and 3D modeling. These innovations have streamlined pre-production workflows for major productions from Marvel Studios and 20th Century Studios.
Digital representations such as avatars, proxies, and digital twins are increasingly being explored in artistic contexts, particularly in relation to identity, presence, and embodiment in virtual environments. The Tate Modern’s film programme Avatars, Proxies and Digital Twins (Feb–May 2025) investigated these themes through curated audiovisual works, offering critical reflections on digital personhood. By engaging with diverse narrative forms, the programme highlighted the sociocultural implications of digital self-representation, prompting discourse on authenticity, agency, and the role of immersive media in shaping future human–machine interaction.